a baloon of mass M is descending at constant acceleration s when a mass m is relased from the ballon it start raising with same acceleration s.assume tht its volume does not change whaty is value of m

Grade:12

2 Answers

Spandan Pratyush, IIT Delhi
13 years ago

Dear Jauneet Singh,

Let the force due to buoyancy be B.

In the first case the equation will be

Mg - B = Ms

In the second case the equation will be

B - (M-m)g = (M-m)s

Since the volume is unchanged, the force of Buoyancy remains unchanged as same amount of air is displaced even now and the new mass is (M-m)

Adding the two equations we get,

mg = (2M-m)s

or m(g+s)=2Ms

or m=2Ms/(g+s)

Please feel free to post as many doubts on our discussion forum as you can.

If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts

Spandan Pratyush

SivaTheGreat
4 years ago
thanks bro 
 

Let the force due to buoyancy be B.

In the first case the equation will be

Mg - B = Ms

In the second case the equation will be

B - (M-m)g = (M-m)s

Since the volume is unchanged, the force of Buoyancy remains unchanged as same amount of air is displaced even now and the new mass is (M-m)

Adding the two equations we get,

mg = (2M-m)s

or m(g+s)=2Ms

or m=2Ms/(g+s)